# On Finding Small Hyper-Gradients in Bilevel Optimization: Hardness Results and Improved Analysis

### v4

@@ Abstract
Bilevel optimization reveals the inner structure of otherwise oblique optimization problems, such as hyperparameter tuning, neural architecture search, and meta-learning. A common goal in bilevel optimization is to minimize a hyper-objective that implicitly depends on the solution set of the lower-level function. Although this hyper-objective approach is widely used, its theoretical properties have not been thoroughly investigated in cases where the lower-level functions lack strong convexity. In this work, we first provide hardness results to show that the goal of finding stationary points of the hyper-objective for nonconvex-convex bilevel optimization can be intractable for zero-respecting algorithms. Then we study a class of tractable nonconvex-nonconvex bilevel problems when the lower-level function satisfies the Polyak-Łojasiewicz (PL) condition. We show a simple first-order algorithm can achieve better complexity bounds of $\tilde{\mathcal{O}}(\epsilon^{-2})$, $\tilde{\mathcal{O}}(\epsilon^{-4})$ and $\tilde{\mathcal{O}}(\epsilon^{-6})$ in the deterministic, partially stochastic, and fully stochastic setting respectively. The complexities in the first two cases are optimal up to logarithmic factors.
{NoHyper}**footnotetext: Equal contributions.footnotetext: Institute for Interdisciplinary Information Sciences, Tsinghua University, China.
@@ 1 Introduction
Bilevel optimization in this form has received increasing attention due to its wide applications in many machine learning problems, including hyperparameter tuning [16, 43], neural architecture search [37, 51, 58, 55], meta-learning [16, 23, 47, 44, 46], out-of-distribution learning [57], adversarial training [18, 50, 32, 33], reinforcement learning [27, 22], causal learning [25, 4].
The hyper-objective approach [10] reformulates Problem (1) by a minimization problem defined below,
@@
Finding stationary points can be especially easy when the lower-level function is strongly convex, because Equation 2 can be simplified to the composite optimization problem below. Specifically, since $Y^{*}(x)$ has only one element when the lower-level function is strongly convex, we have $Y^{*}(x)=\{y^{*}(x)\}$ and
@@
Further, the implicit function theorem [12] implies
@@
[…]
However, due to the prevalence of nonconvex functions in real-world scenarios, the strong convexity assumption may limit the applicability of algorithms. Therefore, our work first aims to whether convexity, as a relaxation of strong convexity, suffices for finding small hyper-gradients efficiently. Specifically, we hope to answer the question below.
@@
We provide a negative answer to the above question. We first prove in Example 3.1 that in the case where $g(x,\cdot)$ is merely convex, $\varphi(x)$ may not have stationary points since $\varphi(x)$ may be discontinuous. Furthermore, we show that the continuity of $\varphi(x)$ can be fully characterized by the Pompeiu–Hausdorff continuity of $Y^{*}(x)$ in Theorem 3.1.
@@
Theorem 1.1 (Informal version of Theorem 3.2).
@@
[…]
Can one achieve (near)-optimal rates for nonconvex-PL bilevel problems with gradient oracles?
[…]
@@
2nd GALET (a) $\tilde{\mathcal{O}}(\kappa^{5}\epsilon^{-2})$ Xiao et al. [52]
[…]
  • (a) Although Xiao et al. [52] did not provide the dependency on $\kappa$ in the complexity, we can calculate by the way in our Remark C.1. Their analysis additionally requires the smallest singular value of $\nabla_{yy}^{2}g(x,y)$ has a constant gap between zero, but such an assumption seems to be redundant by our Remark C.2.
  • (b) We use $p_{1},p_{2},p_{3}$ to denote the polynomial dependency in $\kappa$ since they are not provided by Kwon et al. [29].
Theorem 1.2 (Informal version of Theorem 4.1).
[…]
@@ 2 Preliminaries
This section presents some basic definitions that are commonly used in optimization [42]. To start with, the following definitions describe different orders of smoothness and levels of convexity for a function.
Definition 2.1.
@@
As one relaxation of the above strong convexity condition, the Polyak-Łojasiewicz (PL) condition, independently introduced by Polyak [45] and Lojasiewicz [40], is formally defined as follows.
Definition 2.3.
@@
Compared to the strong convexity, the PL condition allows nonconvexity and multiple minima. The PL condition has wide applications in establishing global convergence of many nonconvex learning problems, including neural network training [7, 35, 21, 31] and optimal control  [14].
@@ 3.1 Stationary Points May not Exist
The following example shows that when the lower-level function only has convexity, $\varphi(x)$ (Equation 2) can be discontinuous and has no stationary points.
Example 3.1 (Lucchetti et al. [41]).
@@ 3.2 Stationary Points May be Intractable to Find
This problem is much simpler because the implicit gradient in Equation 4 disappears as $Y^{*}$ is a fixed set. However, we can show that this problem is hard enough for the following algorithm class.
@@
We define an algorithm class that covers a wide range of existing algorithms designed for bilevel optimization. This definition is inspired by the classical definitions in Nesterov [42] but has an additional structure for bilevel problems (5). We first recall the definition of “zero-respecting” algorithms.
Definition 3.4 (Definition 1 Carmon et al. [6]).
@@
The above zero-respecting algorithm class for bilevel problems subsumes many known algorithms when applied to Problem (5), including:
  1. 1.
[…]
  2. 2.
    many asymptotic algorithms without lower-level strong convexity assumption, such as BGS-Opt [2], BDA [38] and BVFIM [39];
  3. 3.
    the above algorithms with an additional regularizer $\lambda\|y-y_{0}\|^{2}$ for any $\lambda>0$ to $g(y)$.
@@
$f(x,y)=2(x+1)^{2}\sum_{j=q/2}^{q}\psi(y_{[j]}),\quad g(y)=\frac{1}{8}(y_{[1]}-1)^{2}+\frac{1}{8}\sum_{j=1}^{q-1}\left(y_{[j+1]}-y_{[j]}\right)^{2},$
where $\psi(\,\cdot\,):{\mathbb{R}}\rightarrow{\mathbb{R}}$ (defined in Equation 14) is a function with $\psi^{\prime}(0)=\psi(0)=0$. Define the sublevel set $\mathcal{X}:=\{x:\varphi(x)\leq\varphi(0)\}$. There exists numerical constants $c_{1},c_{2}>0$ such that
@@
The idea in our construction is to assign different dimensions in $y$ to functions $\psi$ and $g$ separately and let $g$ be the worst-case function as Nesterov [42] such that a zero-respecting algorithm can only make progress at most one dimension per oracle calls. As the progress in $g$ is slow, it will not affect the dimensions that $\psi$ depends on.
@@
Remark 3.1.
Our results in these two subsections, from two complementary perspectives, motivate us to focus on more well-behaved lower-level functions that: (1) can confer the continuity of $Y^{*}(x)$, to avoid the case as Example 3.1. (2) an algorithm can converge rapidly to a neighborhood of $Y^{*}(x)$, to avoid the case as Theorem 3.2. In the next section, we will show that the PL condition simultaneously satisfies both of these requirements.
@@ 4 Positive Results for Lower-Level Functions Satisfying PL Conditions
As we have shown finding small hyper-gradients of nonconvex-convex bilevel problems is intractable for nonconvex-convex bilevel problems, we turn our attention to the tractable cases. Motivated by recent works [49, 29, 2], we study the case when the lower-level problem satisfies the PL condition.
@@ 4.1 The Assumptions for Nonconvex-PL Bilevel Problems
Without the typical lower-level strong convexity assumption, it is difficult to directly analyze the implicit gradient as Equation 4 since we can not directly use the implicit function theorem. Recently, Kwon et al. [29] proposed a novel way to study the differentiability of $\varphi(x)$ for nonconvex-PL bilevel problems. Instead of directly studying the original hyper-objective $\varphi(x)$, they studied the following value-function penalized hyper-objective as a bridge:
@@
Example 4.1.
Consider Problem (1) with $d_{x}=1$, $d_{y}=2$. Let the upper-level function $f(x,y)=xy_{[1]}$, and the lower-level function $g(x,y)=\frac{1}{2}y_{[2]}^{2}$. The lower-level function is $1$-PL in $y$. Consider the penalized hyper-objective $\varphi_{\sigma}(x)$ in Equation 6. For any $\sigma\geq 0$, if taking minimum in $y$ over all the domain $y\in{\mathbb{R}}^{2}$, $\varphi_{\sigma}(x)$ is not well-defined since $\varphi_{\sigma}(x)=-\infty$ for any $x\neq 0$; if taking minimum in $y$ over a compact domain such as ${\mathcal{Y}}=[0,1]\times[0,1]$, $\varphi_{\sigma}(x)=\min\{x,0\}$ is not differentiable at the point $x=0$.
To guarantee the differentiability of $\varphi_{\sigma}(x)$, Kwon et al. [29] assumed the PL condition not only holds for $g$, but also holds for the penalty function $h_{\sigma}=\sigma f+g$ uniformly for all $\sigma$ near zero. In addition to this assumption, the authors also imposed other standard smoothness assumptions which are typically required in previous works. These assumptions, as stated in Kwon et al. [29], are formally presented below.
Assumption 4.1.
Recall the bilevel problem defined in Equation 1, where $f$ is the upper-level problem, $g$ is the lower-level problem. Let $h_{\sigma}=\sigma f+g$ be the penalty function. Suppose that
@@
Remark 4.1.
[…]
@@
Then we can use Theorem 3.1 to get the continuity of hyper-objective $\varphi(x)$.
[…]
Our starting points are the following lemmas that hold once we have Lipschitz continuity of solution set from Lemma 4.1, which unnecessarily requires strong convexity. Firstly, we can obtain the following result by using the generalized Danskin’s theorem [49] twice, specifically, in both $Y^{*}(x)$ and $Y_{\sigma}^{*}(x)$.
Lemma 4.2 (Shen and Chen [49]).
Recall that $\varphi_{\sigma}(x)$ is the penalized hyper-objective defined in Equation 6. Under Assumption 4.1, $\nabla\varphi_{\sigma}(x)$ exists and takes the form of

### v6

@@ Abstract
Bilevel optimization reveals the inner structure of otherwise oblique optimization problems, such as hyperparameter tuning, neural architecture search, and meta-learning. A common goal in bilevel optimization is to minimize a hyper-objective that implicitly depends on the solution set of the lower-level function. Although this hyper-objective approach is widely used, its theoretical properties have not been thoroughly investigated in cases where the lower-level functions lack strong convexity. In this work, we first provide hardness results to show that the goal of finding stationary points of the hyper-objective for nonconvex-convex bilevel optimization can be intractable for zero-respecting algorithms. Then we study a class of tractable nonconvex-nonconvex bilevel problems when the lower-level function satisfies the Polyak-Łojasiewicz (PL) condition. We show a simple first-order algorithm can achieve complexity bounds of $\tilde{\mathcal{O}}(\epsilon^{-2})$, $\tilde{\mathcal{O}}(\epsilon^{-4})$ and $\tilde{\mathcal{O}}(\epsilon^{-6})$ in the deterministic, partially stochastic, and fully stochastic setting respectively.
keywords:
bilevel optimization, optimization theory, oracle complexity
[…]
@@ 1 Introduction
Bilevel optimization in this form has received increasing attention due to its wide applications in many machine learning problems, including hyperparameter tuning (Franceschi et al., 2018; Pedregosa, 2016), neural architecture search (Liu et al., 2019; Wang et al., 2022; Zoph and Le, 2016; Zhang et al., 2021), meta-learning (Franceschi et al., 2018; Hospedales et al., 2021; Ravi and Larochelle, 2017; Pham et al., 2021; Rajeswaran et al., 2019), out-of-distribution learning (Zhou et al., 2022), adversarial training (Goodfellow et al., 2020; Sinha et al., 2018; Lin et al., 2020a, b), reinforcement learning (Konda and Tsitsiklis, 1999; Hong et al., 2023), causal learning (Jiang and Veitch, 2022; Arjovsky et al., 2019).
The hyper-objective approach (Dempe, 2002) reformulates Problem (1) by a minimization problem defined below,
@@
Finding stationary points can be especially easy when the lower-level function is strongly convex, because Equation (2) can be simplified to the composite optimization problem below. Specifically, since $Y^{*}(x)$ has only one element when the lower-level function is strongly convex, we have $Y^{*}(x)=\{y^{*}(x)\}$ and
@@
Further, the implicit function theorem (Dontchev et al., 2009) implies
@@
This equation enables one to estimate the hyper-gradient $\nabla\varphi(x)$ and perform gradient descent on $\varphi(x)$. AID (Ghadimi and Wang, 2018) and ITD (Ji et al., 2021) estimate $\nabla\varphi(x)$ with Hessian-vector-product oracles. The more recently proposed F2BA (Chen et al., 2023) estimate $\nabla\varphi(x)$ with gradient oracles. All these methods can find a stationary point of $\varphi(x)$.
However, due to the prevalence of nonconvex functions in real-world scenarios, the strong convexity assumption may limit the applicability of algorithms. Therefore, our work first aims to study whether convexity, as a relaxation of strong convexity, suffices for finding small hyper-gradients efficiently. Specifically, we hope to answer the question below.
@@
We provide a negative answer to the above question. We first prove in Example 3.1 that in the case where $g(x,\cdot)$ is merely convex, $\varphi(x)$ may not have stationary points since $\varphi(x)$ may be discontinuous. Furthermore, we show that the continuity of $\varphi(x)$ can be fully characterized by the Pompeiu–Hausdorff continuity of $Y^{*}(x)$ in Theorem 1.
@@
Theorem 1.1 (Informal version of Theorem 2)
@@
The PL condition can pose nontrivial challenges since we have neither Equation (3) nor (4) in this case. Researchers have provided novel analyses in this case. Xiao et al. (2023) proposed a Hessian-vector-product-based method GALET with non-asymptotic convergence to KKT points of the gradient-based reformulated problem (Equation (11)) when $g$ is PL in $y$. Kwon et al. (2024b) proved the differentiability of $\varphi(x)$ when the penalty function $\sigma f+g$ is uniformly PL in $y$ for all $\sigma$ in the neighborhood of zero, and showed a proximal variant of F2BA, which we call Prox-F2BA, can find a stationary point of $\varphi(x)$. Based on their differentiability result, we show that GALET also converges to a stationary point of $\varphi(x)$ in Appendix B.
Although Prox-F2BA (Kwon et al., 2024b) has been shown to converge to an $\epsilon$-stationary point of $\varphi(x)$ with $\tilde{\mathcal{O}}(\epsilon^{-3})$ first-order oracle calls, the rate is worse than the ${\mathcal{O}}(\epsilon^{-2})$ rate of gradient descent on nonconvex single-level optimization. Therefore, listed as an important future direction, Kwon et al. (2024b) asked the question below.
Can one achieve the ${\mathcal{O}}(\epsilon^{-2})$ rate for nonconvex-PL bilevel problems with gradient oracles?
We give a positive answer to this question. Interestingly, we prove that the original F2BA (Chen et al., 2023) can already achieve this fast convergence rate. Our improvement over Kwon et al. (2024b) comes from establishing a tighter bound on the smoothness constant of $\varphi(x)$.
@@
2nd GALET (a) $\tilde{\mathcal{O}}(\kappa^{5}\epsilon^{-2})$ Xiao et al. (2023)
1st Prox-F2BA (b) $\tilde{\mathcal{O}}(\kappa^{p_{1}}\epsilon^{-3})$ $\tilde{\mathcal{O}}(\kappa^{p_{2}}\epsilon^{-5})$ $\tilde{\mathcal{O}}(\kappa^{p_{3}}\epsilon^{-7})$ Kwon et al. (2024b)
1st F2BA $\tilde{\mathcal{O}}(\kappa^{4}\epsilon^{-2})$ $\tilde{\mathcal{O}}(\kappa^{5}\epsilon^{-4})$ $\tilde{\mathcal{O}}(\kappa^{11}\epsilon^{-6})$ This Paper
  • (a) Although Xiao et al. (2023) did not provide the dependency on $\kappa$ in the complexity, we can calculate by the way in our Remark B.1. Their analysis additionally requires the smallest singular value of $\nabla_{yy}^{2}g(x,y)$ has a constant gap between zero, but such assumption seems to be redundant by our Remark B.2.
  • (b) We use $p_{1},p_{2},p_{3}$ to denote the polynomial dependency in $\kappa$ since they are not provided by Kwon et al. (2024b).
Theorem 1.2 (Informal version of Theorem 6)
Under regular conditions as in Kwon et al. (2024b), F2BA can provably find an $\epsilon$-stationary point of $\varphi(x)$ with $\tilde{\mathcal{O}}(\epsilon^{-2})$ first-order oracle calls for nonconvex-PL bilevel problems.
We also extend our analysis to stochastic bilevel problems when an algorithm only has access to a noisy estimator of $\nabla f$ and $\nabla g$. Under the partially stochastic setting when the noise is only in $\nabla g$, we prove the stochastic F2BA has the $\tilde{\mathcal{O}}(\epsilon^{-4})$ first-order oracle complexity, matching SGD for single-level nonconvex minimization (Arjevani et al., 2023). Under the more general fully stochastic setting when noise appears both in $\nabla f$ and $\nabla g$, we prove the method has an $\tilde{\mathcal{O}}(\epsilon^{-6})$ complexity.
Compared with Xiao et al. (2023), our deterministic F2BA achieves the same $\tilde{\mathcal{O}}(\epsilon^{-2})$ rate without the assistance of Hessian-vector-product oracles. Our method also has a better dependency on $\kappa$ since we do not use a squared trick on the Hessian of $g$. And we additionally study stochastic problems that have not been studied by Xiao et al. (2023). Compared with Kwon et al. (2024b), we strictly improve the complexities for both deterministic and stochastic cases under the same assumptions. We compare our results with prior works in Table 1, and leave a more detailed introduction of related works in Appendix A.
@@ 2 Preliminaries
This section presents some basic definitions that are commonly used in optimization (Nesterov, 2018). To start with, the following definitions describe different orders of smoothness and levels of convexity for a function.
Definition 2.1
@@
As one relaxation of the above strong convexity condition, the Polyak-Łojasiewicz (PL) condition, independently introduced by Polyak (1967) and Lojasiewicz (1963), is formally defined as follows.
Definition 2.3
@@
Compared to the strong convexity, the PL condition allows nonconvexity and multiple minima. The PL condition has wide applications in establishing global convergence of many nonconvex learning problems, including neural network training (Charles and Papailiopoulos, 2018; Liu et al., 2022b; Hardt and Ma, 2016; Li et al., 2018) and optimal control  (Fazel et al., 2018).
@@ 3.1 Stationary Points May not Exist
The following example shows that when the lower-level function only has convexity, $\varphi(x)$ (Equation (2)) can be discontinuous and has no stationary points.
Example 3.1 (Lucchetti et al. (1987))
@@
Remark 3.1.
In the above example, the lower-level problem is a constrained optimization in $y$. We can also give a similar counter-example for unconstrained problems by replacing $I_{C}(y)$ with a smoothed surrogate $h(y)=(y-1){\mathbb{I}}[y\geq 1]-y{\mathbb{I}}[y\leq 0]$ and then letting $g(x,y)=xy+h(y)$.
@@ 3.2 Stationary Points May be Intractable to Find
This problem is much simpler because the implicit gradient in Equation (4) disappears as $Y^{*}$ is a fixed set. However, we can show that this problem is hard enough for the following algorithm class.
@@
We define an algorithm class that covers a wide range of existing algorithms designed for bilevel optimization. This definition is inspired by the classical definitions in Nesterov (2018) but has an additional structure for bilevel problems (5). We first recall the definition of “zero-respecting” algorithms.
Definition 3.4 (Definition 1 Carmon et al. (2021))
@@
The above zero-respecting algorithm class for bilevel problems subsumes many known algorithms when applied to Problem (5), including: AID (Ghadimi and Wang, 2018), ITD (Ji et al., 2021), GALET (Xiao et al., 2023), (Prox)-F2BA (Kwon et al., 2024b, 2023; Chen et al., 2023), FdeHBO (Yang et al., 2024), BGS-Opt (Arbel and Mairal, 2022), BDA (Liu et al., 2020), BVFIM (Liu et al., 2021a), PDBO (Sow et al., 2022), SLM (Lu, 2023), LV-HBA (Yao et al., 2024).
@@
$f(x,y)=2(x+1)^{2}\sum_{j=q/2}^{q}\psi(y_{[j]}),\quad g(y)=\frac{1}{8}(y_{[1]}-1/\sqrt{q})^{2}+\frac{1}{8}\sum_{j=1}^{q-1}\left(y_{[j+1]}-y_{[j]}\right)^{2},$
where $\psi(\,\cdot\,):{\mathbb{R}}\rightarrow{\mathbb{R}}$ (defined in Equation (14)) is a function with $\psi^{\prime}(0)=\psi(0)=0$. Define the sublevel set $\mathcal{X}:=\{x:\varphi(x)\leq\varphi(0)\}$. There exists numerical constants $c_{1},c_{2}>0$ such that
@@
In our construction, we let $\psi(t)$ be a function which is $\frac{1}{2}t^{2}$ near zero, while remaining bounded when $|t|$ is large. We assign different dimensions in $y$ to functions $\psi$ and $g$ separately. Note that $g$ is designed such that any zero-respecting algorithm can only make progress at most one dimension per oracle calls. As the progress in $g$ is slow, it will not affect the dimensions that $\psi$ depends on.
@@
Remark 3.1
Our results in these two subsections, from two complementary perspectives, motivate us to focus on more well-behaved lower-level functions that: (1) can confer the continuity of $Y^{*}(x)$, to avoid the case as Example 3.1. (2) an algorithm can converge rapidly to a neighborhood of $Y^{*}(x)$, to avoid the case as Theorem 2. In the next section, we will show that the PL condition simultaneously satisfies both of these requirements.
@@ 4 Positive Results for Lower-Level Functions Satisfying PL Conditions
As we have shown finding small hyper-gradients of nonconvex-convex bilevel problems is intractable for nonconvex-convex bilevel problems, we turn our attention to the tractable cases. Motivated by recent works (Shen and Chen, 2023; Kwon et al., 2024b; Arbel and Mairal, 2022), we study the case when the lower-level problem satisfies the PL condition.
@@ 4.1 The Assumptions for Nonconvex-PL Bilevel Problems
Without the typical lower-level strong convexity assumption, it is difficult to directly analyze the implicit gradient as Equation (4) since we can not directly use the implicit function theorem. Recently, Kwon et al. (2024b) proposed a novel way to study the differentiability of $\varphi(x)$ for nonconvex-PL bilevel problems. Instead of directly studying the original hyper-objective $\varphi(x)$, they studied the following value-function penalized hyper-objective as a bridge:
@@
Example 1.
Consider Problem (1) with $d_{x}=1$, $d_{y}=2$. Let the upper-level function $f(x,y)=xy_{[1]}$, and the lower-level function $g(x,y)=\frac{1}{2}y_{[2]}^{2}$. The lower-level function is $1$-PL in $y$. Consider the penalized hyper-objective $\varphi_{\sigma}(x)$ in Equation (6). For any $\sigma\geq 0$, if taking minimum in $y$ over all the domain $y\in{\mathbb{R}}^{2}$, $\varphi_{\sigma}(x)$ is not well-defined since $\varphi_{\sigma}(x)=-\infty$ for any $x\neq 0$; if taking minimum in $y$ over a compact domain such as ${\mathcal{Y}}=[0,1]\times[0,1]$, $\varphi_{\sigma}(x)=\min\{x,0\}$ is not differentiable at the point $x=0$.
To guarantee the differentiability of $\varphi_{\sigma}(x)$, Kwon et al. (2024b) assumed the PL condition not only holds for $g$, but also holds for the penalty function $h_{\sigma}=\sigma f+g$ uniformly for all $\sigma$ near zero. In addition to this assumption, the authors also imposed other standard smoothness assumptions which are typically required in previous works. These assumptions, as stated in Kwon et al. (2024b), are formally presented below.
Assumption 4.1
Recall the bilevel problem defined in Equation (1), where $f$ is the upper-level problem, $g$ is the lower-level problem. Let $h_{\sigma}=\sigma f+g$ be the penalty function. Suppose that
@@
Remark 4.1
Note that all the assumptions are the same as Kwon et al. (2024b), except (a) may seem different from the Prox-EB assumption (Assumption 1 in Kwon et al. (2024b)). However, in Appendix C we show these two assumptions can imply each other. Our narrative uses the PL condition and is therefore more convenient. As a by-product, we show the proximal operator in Kwon et al. (2024b) is unnecessary, and the original F2BA (Chen et al., 2023) can also converge under PL conditions.
@@
Then we can use Theorem 1 to get the continuity of hyper-objective $\varphi(x)$.
4.2 F2BA Can also be Applied to Nonconvex-PL Bilevel Problems
Although F2BA is originally proposed for nonconvex-strongly-convex bilevel problems, we show that F2BA can also be applied to nonconvex-PL bilevel problems in this section.
Our starting points are the following lemmas that hold once we have Lipschitz continuity of solution set from Lemma 2, which unnecessarily requires strong convexity. Firstly, we can obtain the following result by using the generalized Danskin’s theorem (Shen and Chen, 2023) twice, specifically, in both $Y^{*}(x)$ and $Y_{\sigma}^{*}(x)$.
Lemma 3 (Shen and Chen (2023)).
Recall that $\varphi_{\sigma}(x)$ is the penalized hyper-objective defined in Equation (6). Under Assumption 4.1, $\nabla\varphi_{\sigma}(x)$ exists and takes the form of